OpenStudy (anonymous):

What is the sum of the arithmetic sequence 8, 14, 20 …, if there are 22 terms?

OpenStudy (anonymous):

it seems the secquences is x+6

OpenStudy (anonymous):

so with that being said keep adding 6 until you get to the 22th term then just add them together

OpenStudy (dls):

Where, a=8 d=common difference n=number of terms=22

OpenStudy (cwrw238):

DLS r u sure thats correct formula - isn't n/2 not a/2

OpenStudy (dls):

Sorry,if its n/2,doing after a long time Sn=n/2(2a+(n-1)d Sn=1(16+21*6) Sn=11(142) Sn=1562

OpenStudy (cwrw238):

yes i just checked - so its 11(16 + 21*6)

OpenStudy (dls):

wth is going on sorry its Sn=11 (16+21*6)* not one ! (2nd line)